# End-to-End Portfolio Optimization with Quantum Annealing

### v1

@@ Abstract
With rapid technological progress reshaping the financial industry, quantum technology plays a critical role in advancing risk management, asset allocation, and financial strategies. Realizing its full potential requires overcoming challenges like quantum hardware limits, algorithmic stability, and implementation barriers. This research explores integrating quantum annealing with portfolio optimization, highlighting quantum methods’ ability to enhance investment strategy efficiency and speed. Using hybrid quantum-classical models, the study shows combined approaches effectively handle complex optimization better than classical methods. Empirical results demonstrate a portfolio increase of 200,000 Indian Rupees over the benchmark. Additionally, using rebalancing leads to a portfolio that also surpasses the benchmark value.
Index Terms:
Quantum annealing, Asset selection, Asset allocation, QUBO, Combinatorial optimization.
@@ I Introduction
Portfolio optimization is a core financial task that seeks to allocate assets in a manner that maximizes returns while minimizing risk. Traditionally, this process relies on mathematical models such as Modern Portfolio Theory (MPT) and Mean-Variance Optimization (MVO), which use classical algorithms to assess and predict the behavior of various asset classes. However, these classical solutions often encounter limitations in scalability and speed, particularly when dealing with large datasets or complex asset interactions.
@@
This paper outlines these methodologies and provides both empirical analysis and theoretical insights into their effectiveness. By integrating quantum annealing into portfolio optimization, we aim to establish a more adaptive and efficient framework that can operate in the dynamic and often unpredictable realm of financial markets.
@@ II Background and Related Work
Quantum annealing (QA) [1] has been of primary interest since the early applications of quantum technology in finance, D-Wave Systems has been developing quantum annealing hardware or quantum annealers [2], [3], [4] for over two decades. Their product line includes the 2000QTM, which features 2,000 qubits, alongside more recent models in the Advantage family. The latest of these, the Advantage2 QPU, contains over 1,200 qubits and offers a 20-way connectivity. According to D-Wave, this architecture delivers better solutions 87% of the time compared to the prior Advantage system[5]. Rosenberg et al. [6] and Glover et al. [7] were among the first to formulate optimization problems into Quadratic Unconstrained Binary Optimization (QUBO) problems and showed they has better feasibility for quantum annealers. Their approach effectively demonstrated how complex combinatorial problems in finance could be mapped onto quantum architectures. There has also been a new formulation called polynomial unconstrained binary optimization (PUBO), where the penalty terms take binary values if the constraints are linear [8].
Venturelli and Kondratyev [9] applied quantum annealing techniques to multi-period portfolio optimization problems, adding transaction costs and proving actual scalability using genuine quantum annealers. Their trials revealed not only computing gains, but also practical elements applicable to actual trading settings. The primary motivation behind this work is from [10], [11], [12].
Significant research has been carried out using gate-based quantum algorithms, particularly the Quantum Approximate Optimization Algorithm (QAOA) developed by [13], [14] and [15] employed QAOA for portfolio optimization and reported improvements over traditional approaches in simulated environments. However, these results do not constitute a rigorous proof that QAOA outperforms classical methods universally. Their findings suggest that, while QAOA is still in its early stages, it may offer practical advantages as quantum technology advances.
Brandhofer et al. [16], [17], [18] conducted benchmark experiments to compare the performance of the QAOA and quantum annealing against classical heuristic methods such as simulated annealing and genetic algorithms. In these experiments, the authors evaluated the methods on specific combinatorial optimization problems by measuring factors like solution quality and computational efficiency. Their analysis indicates that current quantum approaches are limited by hardware constraints—including qubit count, noise levels, and connectivity issues—which restrict their performance relative to established classical algorithms. The study further suggests that improvements in quantum hardware and error correction techniques could lead to performance enhancements in quantum optimization methods, enabling them to be more competitive in future applications. This research [19] applies the Variational Quantum Eigensolver (VQE) to portfolio optimization, proposing hardware-efficient ansatzes suitable for near-term quantum devices.
Recent comparative studies like [20] have further validated the practical advantage of hybrid quantum-classical models using real-world portfolio constraints and benchmarks, reinforcing the necessity of hybrid approaches under current hardware limitations. Acharya et al. [21] proposed a decomposition pipeline that partitions large portfolio optimization problems into manageable subproblems, enabling the application of near-term quantum devices for solving complex financial models. The range and applicability of quantum solutions in portfolio optimization have been greatly expanded by recent developments in hybrid quantum-classical approaches. Hybrid methods refer to algorithms that combine quantum techniques with classical optimization strategies. In these approaches, the quantum component is typically used for parts of the problem where quantum properties, such as superposition and entanglement, can be advantageous, while classical algorithms address the remaining computational tasks.
In order to effectively utilize quantum advantages, Bouland et al. [22] emphasized hybrid algorithms as a feasible intermediate step that combines quantum calculations with classical optimization techniques. As demonstrated by Egger et al. [23], who created useful hybrid quantum algorithms especially suited for realistic financial optimization challenges, this strategy has garnered a lot of interest.
Quantum approaches have also been thoroughly investigated in other financial applications, such as risk management, option pricing, and arbitrage detection, in addition to portfolio optimization. To explain how quantum amplitude estimation can greatly speed up value-at-risk (VaR) and conditional VaR computations, which are essential for financial risk management, Woerner and Egger [24] showed. The practical application of quantum algorithms for option pricing was also shown by Stamatopoulos et al. [25], who showed computational complexity speedups over conventional Monte Carlo simulations.
Additionally, Orús et al. [26] offered a thorough analysis of quantum applications in the finance industry, highlighting portfolio optimization as a crucial field where quantum techniques might be quite advantageous. In addition to highlighting possible computing benefits, their evaluation pointed out existing drawbacks such coherence times, scalability, and quantum noise, setting the stage for further study.
@@ III Methodology
where $\tau$ represents the entire annealing time. According to the adiabatic theorem, if this evolution is slow enough, the system will remain in its instantaneous ground state throughout, resulting in the ground state of $H_{1}$ when $t=\tau$. Typically, the initial Hamiltonian is defined as $H_{0}=-\sum_{i}\sigma^{x}_{i}$, where $\sigma^{x}_{i}$ represents the Pauli-X operator acting on the $i$-th qubit. The problem Hamiltonian takes the form of an Ising model:
$H_{1}=\sum_{i>j}J_{ij}\sigma^{z}_{i}\sigma^{z}_{j}+\sum_{i}h_{i}\sigma^{z}_{i},$ (2)
where $J_{ij}$ represents couplings between qubits $i$ and $j$, $h_{i}$ denotes external fields, and $\sigma^{z}_{i}$ is the Pauli-Z operator.
@@ III-A Modern Portfolio Theory (MPT) formulation
The number of shares $x_{i}$ affects the portfolio risk through its contribution to the overall variance term. Since the variance is calculated using the covariances $\sigma_{ij}$, the individual standard deviation $\sigma^{z}_{i}=\sqrt{\sigma_{ii}}$ indirectly influences the portfolio’s total risk depending on the amount invested in asset $i$, i.e., $p_{i}x_{i}$. Thus, higher allocations to more volatile stocks (with high $\sigma^{z}_{i}$) increases the portfolio variance.
@@ III-B Sharpe ratio maximization
Sharpe $\text{Sharpe ratio}=\frac{E[R_{p}-R_{f}]}{\sigma_{p}}=\frac{E[R_{p}-R_{f}]}{\sqrt{var[R_{p}-R_{f}]}},$ (5)
@@ III-C Mean-variance portfolio optimization (MVO)
We need to remove the constraint and formulate the problem as a QUBO to make it suitable for quantum computer. The equality constraint $1^{T}x=B$ is mapped to a penalty term $(1^{T}x-B)^{2}$ which is scaled by a parameter (Lagrange multiplier) and subtracted from the objective function.
$\min_{x\in\{0,1\}^{n}}qx^{T}Cx-\mu^{T}x-\lambda(1^{T}x-B)^{2},$ (8)
@@
  • Budget constraint penalty: $-\lambda(1^{T}x-B)^{2}$
@@
We convert the QUBO problem (8) into an Ising Hamiltonian [30]. As the Ising model has spin variables $s_{i}\in\{-1,1\}$, ${x_{i}}$ is transformed to (1+${s_{i}}$)/2. The classical Ising Hamiltonian of our QUBO problem is given as follows:
@@
[…]
where ${J_{ij}}$ is the coupling term between two variables. To implement this Hamiltonian on a D-Wave quantum annealer, we first transform the problem into the language of the Pauli $Z$ operators, which have eigenvalues of ±$\pm 1$. The Ising Hamiltonian is:
@@ III-E2 Portfolio Health check
III-E3 Buy Idea
For identified risky companies (say N number of risky companies), all their stocks are sold as of now. The sales revenue generated is used as the new budget to purchase stocks again.
  • Identify the sectors of the companies whose stocks are sold
  • Obtain the list of companies in those sectors excluding the ones whose stocks are sold.
  • Perform portfolio optimization using the list of companies and the new budget. The output of the algorithm is a list of $N$ companies along with the number of stocks to purchase of each company.
  • Buy those stocks and update the holdings.
Fig. 2 illustrates the flow digarm of rebalancing, where fully quantum approach has been used in generating the results.
@@ IV Results and Discussion
TABLE I: Attribution Report of a real world portfolio. Note: Port = Portfolio, Bench = Benchmark, Ret = Return. Values represent percentage weights and returns.
@@
Table II presents a comparative analysis of the asset allocation weights determined by our algorithm against those provided by the bank’s benchmark. This table also includes key financial metrics such as risk, return, Sharpe ratio, and diversification ratios. Our algorithm’s allocation strategy results in superior financial performance. Specifically, the algorithm’s weight distribution yields higher returns and a more favorable Sharpe ratio in comparison to the benchmark distribution. This demonstrates the efficacy of our method in optimizing asset allocation to achieve enhanced performance metrics.
@@
In Fig. 3, the red line represents the portfolio value of the benchmark data for our analysis. Conversely, the blue line depicts the portfolio value achieved by our algorithm. Initially, both the user’s and the algorithm’s portfolios were valued at 1.6 million Indian Rupees. However, after a period of 13 months, the individual investor’s portfolio appreciated to approximately 1.95 million Indian Rupees. In contrast, the algorithm’s portfolio increased to 2.2 million Indian Rupees. This illustrates a notable difference, with the algorithm’s portfolio outperforming the user’s by approximately 0.3 million Indian Rupees.

### v2

# End-to-End Portfolio Optimization with Hybrid Quantum Annealing

@@ Abstract
Hybrid quantum-classical optimization has emerged as a promising direction for addressing financial decision problems under current quantum hardware constraints. In this work we present a practical end-to-end portfolio optimization pipeline that combines (i) a continuous mean-variance and Sharpe-ratio formulation, (ii) a QUBO/CQM-based discrete asset selection stage solved using D-Wave’s hybrid quantum annealing solver, (iii) classical convex optimization for computing optimal asset weights, and (iv) a quarterly rebalancing mechanism. Rather than claiming quantum advantage, our goal is to evaluate the feasibility and integration of these components within a deployable financial workflow. We empirically compare our hybrid pipeline against a fund manager in real time and indexes used in Indian stock market. The results indicate that the proposed framework can construct diversified portfolios and achieve competitive returns. We also report computational considerations and scalability observations drawn from the hybrid solver behaviour. While the experiments are limited to moderate sized portfolios dictated by current annealing hardware and QUBO embedding constraints, the study illustrates how quantum assisted selection and classical allocation can be combined coherently in a real-world setting. This work emphasizes methodological reproducibility and practical applicability, and aims to serve as a step toward larger-scale financial optimization workflows as quantum annealers continue to mature.
@@ I Introduction
Portfolio optimization is a core financial task that seeks to allocate assets in a manner that maximizes returns while minimizing risk. Traditionally, this process relies on mathematical models such as Modern Portfolio Theory (MPT)[1] and Mean-Variance Optimization (MVO)[2], which use classical algorithms to assess and predict the behavior of various asset classes. However, these classical solutions often encounter limitations in scalability and speed[3], [4], [5], [6], particularly when dealing with large datasets or complex asset interactions.
@@
This paper outlines these methodologies and provides both empirical analysis and theoretical insights into their effectiveness. By integrating quantum annealing into portfolio optimization [7, 8], we aim to establish a more adaptive and efficient framework that can operate in the dynamic and often unpredictable realm of financial markets.
@@
This study does not claim that quantum annealing or the hybrid CQM solver provides superior runtime or scalability relative to tuned classical solvers. We refrain from drawing conclusions about asymptotic scaling or computational speed. The contribution of this work lies in demonstrating a feasible, end-to-end integration of quantum-assisted discrete selection with classical allocation and rebalancing on real financial data, rather than establishing quantum computational advantage.
@@ II Background and Related Work
Quantum annealing (QA) [9] has emerged as a practical approach for solving combinatorial optimization problems in finance. D-Wave Systems has developed several generations of quantum annealers, including the 2000QTM and the Advantage family, which employ transverse-field Ising architectures for energy minimization [10, 11, 12, 13]. These systems implement optimization problems formulated as Quadratic Unconstrained Binary Optimization (QUBO) or Ising models, following the early mappings proposed by Rosenberg [14] and Glover [8]. Recent work has also extended this framework to polynomial unconstrained binary optimization (PUBO) [15].
Applications of QA to financial modeling began with Venturelli and Kondratyev [16], who addressed multi-period portfolio optimization with transaction costs on physical D-Wave hardware, demonstrating early scalability. Subsequent studies have adopted both annealing and gate-based paradigms. The Quantum Approximate Optimization Algorithm (QAOA) [17] has been explored for portfolio allocation and rebalancing [18, 19], though results remain hardware-limited. Comparative benchmarks [20, 21, 22] indicate that quantum methods currently perform comparably to classical heuristics such as simulated annealing and genetic algorithms, with hardware precision and connectivity constraints being the main bottlenecks. Hybrid strategies that combine classical optimizers with quantum subroutines are therefore viewed as the most feasible near-term route [23]. Related advances include decomposition-based hybrid frameworks for large portfolios [24] and hybrid eigensolver models [25].
Beyond portfolio optimization, quantum techniques have been applied to risk analysis, option pricing, and arbitrage detection. For instance, Woerner and Egger [26] employed amplitude estimation to accelerate Value-at-Risk (VaR) and conditional VaR computations, while Stamatopoulos et al. [27] demonstrated quantum-based option pricing speedups. Orús et al. [28] provided a comprehensive survey of such applications, emphasizing both the promise and the practical limitations of current devices.
While several prior works have explored hybrid quantum–classical approaches for portfolio optimization [16, 29, 30], most studies remain constrained by simplified datasets, limited covariance structures, and a lack of dynamic rebalancing. Few integrate the entire pipeline from mean–variance formulation through QUBO mapping, annealing, decoding, and time-series-based rebalancing within a unified financial framework. Our work addresses these gaps by presenting a comprehensive end-to-end quantum annealing model that (1) reformulates continuous mean–variance and Sharpe-ratio objectives into QUBO form without ad-hoc linearization, (2) employs D-Wave’s hybrid solver to capture real-market correlations from the Indian equity market, and (3) demonstrates reproducible rebalancing that outperforms the fund manager in both return and risk metrics.
@@ III Methodology
where $\tau$ represents the entire annealing time. According to the adiabatic theorem, if this evolution is slow enough, the system will remain in its instantaneous ground state throughout, resulting in the ground state of $H_{1}$ when $t=\tau$. Typically, the initial Hamiltonian is defined as $H_{0}=-\sum_{i}X_{i}$, where $X_{i}$ represents the Pauli-X operator acting on the $i$-th qubit. The problem Hamiltonian takes the form of an Ising model:
$H_{1}=\sum_{i>j}J_{ij}Z_{i}Z_{j}+\sum_{i}h_{i}Z_{i},$ (2)
where $J_{ij}$ represents couplings between qubits $i$ and $j$, $h_{i}$ denotes external fields, and $Z_{i}$ is the Pauli-Z operator.
@@ III-A Modern Portfolio Theory (MPT) formulation
The number of shares $x_{i}$ affects the portfolio risk through its contribution to the overall variance term. Since the variance is calculated using the covariances $\sigma_{ij}$, the individual standard deviation $\sigma_{i}=\sqrt{\sigma_{ii}}$ indirectly influences the portfolio’s total risk depending on the amount invested in asset $i$, i.e., $p_{i}x_{i}$. Thus, higher allocations to more volatile stocks (with high $\sigma_{i}$) increases the portfolio variance.
@@ III-B Sharpe ratio maximization
$\text{Sharpe ratio}=\frac{E[R_{p}-R_{f}]}{\sigma_{p}}=\frac{E[R_{p}-R_{f}]}{\sqrt{var[R_{p}-R_{f}]}},$ (5)
@@ III-C Mean-variance portfolio optimization (MVO)
cov for covarience but not sigma as it is confusing To convert the budget equality constraint $1^{\top}x=B$ into an unconstrained QUBO, we add a positive quadratic penalty term $\lambda_{b}(1^{\top}x-B)^{2}$ to the objective. For a minimization problem, violations of the constraint must increase the objective, hence the penalty enters with a positive sign.
$\min_{x\in\{0,1\}^{n}}\;q\,x^{\top}Cx\;-\;\mu^{\top}x\;+\;\lambda_{b}\,(1^{\top}x-B)^{2},$ (8)
@@
  • Budget constraint penalty: $\lambda_{b}(1^{T}x-B)^{2}$
@@
[…]
@@
$\quad\text{s.t.}\quad s_{i}\in\{-1,1\}\quad\forall i,$
where $J_{ij}$} is the coupling term between two variables. To implement this Hamiltonian on a D-Wave quantum annealer, we first transform the problem into the language of the Pauli $Z$ operators, which have eigenvalues of $\pm 1$. The Ising Hamiltonian is:
@@ III-E2 Portfolio Health check
A detailed description of rebalancing technique is given below:
Input: initial holdings $H^{(0)}$; sector map $s(i)$; budget $B$; risk-aversion $q$; sell count $K_{\text{sell}}$.
Output: new portfolio, its portfolio value, and risk/return metrics. (1) Estimate returns/covariance per quarter: compute HP-filtered prices, log-returns on, then annualized mean $\mu^{(k)}$ and covariance $\Sigma^{(k)}$. (2) Identify underperformers: rank the current holdings in $H^{(0)}$ by $\mu^{(k)}$; mark the lowest as underperformers and liquidate them. (3) Sector-matched candidate set: collect replacements from the same sectors as the sold names, available in the universe at quarter end. (4) Discrete selection (QUBO): solve equation 8 using a classical SA sampler for the QUBO to pick new assets. (5) Integer allocation (CQM, hybrid): for chosen names, solve a single-period CQM on D-Wave’s hybrid solver, update holdings and wallet. (6) Record metrics: compute portfolio value, return, risk, and Sharpe ratio of the new portoflio. In our implementation, “risky companies” refers not to high-variance names but to underperformers by expected return $\mu$.
Fig. 2 illustrates the flow diagram of rebalancing, where fully quantum approach has been used in generating the results.
@@
Although the underlying optimization models follow the standard Modern Portfolio Theory and Mean–Variance formulations, our implementation introduces several key innovations in translating these into a QUBO/Ising framework. First, the continuous portfolio weights are discretized through a structured mapping that preserves convexity and avoids arbitrary thresholding. Second, an adaptive penalty factor $\lambda$ is applied within the constraint Hamiltonian $H_{c}$ to dynamically balance the trade-off between the return–risk objective and the budget or cardinality constraints during the annealing process. Third, after sampling from the annealer, a normalization-based decoding step ensures that the resulting bitstrings correspond to feasible weight allocations whose total sum equals the capital budget. Finally, the Hamiltonian includes a temporal continuity term that links portfolios across rebalancing intervals, encouraging smoother transitions and reduced turnover. These additions collectively form the methodological contribution of our work and distinguish the proposed framework from existing static or single-period QUBO formulations used in earlier quantum finance studies.
@@ IV Results and Discussion
TABLE I: Attribution Report of a real world portfolio. Note: Port = Portfolio, Bench = Benchmark, Ret = Return. Start date = 1 Januray 2023, end date = 29 February 2024 Values represent percentage weights and returns.
@@
Table II presents a comparative analysis of the asset allocation weights determined by our algorithm against those provided by the bank’s benchmark. This table also includes key financial metrics such as risk, return, Sharpe ratio, and diversification ratios. Our algorithm’s allocation strategy results in superior financial performance. Specifically, the algorithm’s weight distribution yields higher returns and a more favorable Sharpe ratio in comparison to the benchmark distribution. This demonstrates the efficacy of our method in optimizing asset allocation to achieve enhanced performance metrics. It is important to note that the weight allocations in Table II are derived from the classical MIQP optimizer, while the quantum solver is utilized only during the rebalancing stage. The overweighting of ICICI Bank, ITC, and Bharti Airtel reflects their superior expected return-to-risk characteristics within the classical covariance structure.
@@
Algo 1 summarizes the complete end-to-end workflow implemented in this study. The pipeline begins with data ingestion and preprocessing, where daily prices are converted into log-returns and used to estimate the expected-return vector and covariance matrix for each rebalancing window. These estimates feed into a discrete asset-selection module, formulated as a QUBO and solved using a quantum-assisted (or simulated) annealing routine. The selected assets are then passed to an integer-constrained CQM allocator that computes feasible share allocations under budget, sector, and cardinality constraints. Finally, a quarterly rebalancing mechanism identifies underperforming holdings, replaces them through a repeat of the selection stage, and updates the portfolio’s evaluation metrics. This unified structure provides a transparent hybrid framework that integrates quantum-assisted selection with classical optimization and realistic trading constraints.
Algorithm 1 End-to-End Portfolio Optimization and Rebalancing via Quantum Annealing
1:Input: Historical prices $P_{i,t}$ for assets $\mathcal{U}$; rebalancing interval $\Delta t=quaterly$; constants $q,\lambda,K_{\text{sell}},B$
2:Output: Optimized quarterly portfolios $\{w^{(k)}\}$ with risk–return metrics
3:
4:Data Preparation
5:Retrieve adjusted close prices for $\mathcal{U}$ from NSE/BSE between $t_{\text{start}}$ and $t_{\text{end}}$
6:Compute log-returns, annualized mean $\mu$, and covariance matrix $\Sigma$
7:Apply HP filter ($\lambda_{\text{HP}}=6.25$) to smooth price trends
8:
9:Optimization Stage
10:for each rebalancing window $k$ do
11:  Formulate mean–variance objective: $\min_{x}\;q\,x^{\top}\Sigma x-\mu^{\top}x$
12:  Encode budget & cardinality constraints: $H=H_{\text{obj}}+\lambda H_{c}$
13:  Convert $H$ to QUBO using QuadraticProgramToQubo
14:  Solve using SimulatedAnnealingSampler (num_reads = 5000)
15:  Decode lowest-energy bitstring and normalize to obtain $w^{(k)}$
16:
17:Rebalancing and Evaluation
18:for each quarter $k\rightarrow k{+}1$ do
19:  Rank current holdings by expected return $\mu$
20:  Sell the $K_{\text{sell}}=4$ lowest-return assets to free capital
21:  Identify candidate replacements from the same sectors
22:  Re-solve the QUBO on this subset to select $K_{\text{sell}}$ new assets
23:  Allocate integer shares
24:  Update holdings and compute portfolio value, risk, return, and Sharpe ratio
In Fig. 3, the red line represents the portfolio value of the benchmark data for our analysis. Conversely, the blue line depicts the portfolio value achieved by our algorithm. Over the 13-month period, both the individual investor and the algorithm started with the same portfolio value of 1.6 million Indian Rupees. By the end of this period, the investor’s portfolio had grown to approximately 1.95 million, reflecting an absolute gain of 0.35 million rupees, or about 21.9% growth. In comparison, the algorithm’s portfolio reached 2.2 million rupees, marking a larger increase of 0.6 million, or 37.5%. This means the algorithm not only delivered a higher return overall but also outperformed the investor by around 0.25 million rupees in absolute terms, which translates to a performance gap of roughly 15.6 percentage points.
@@
We did not restrain from testing the above methods with one single portfolio. Below are more results on different portfolios with various benchmarks used in the Indian market. In fig. 5 we apply the rebalancing technique to a portfolio by a real user which consists of 70 assets, where the user bought 100 shares of each asset. He started his portfolio on the 1st of September 2023 and stopped all his transactions on 25th September 2024. The benchmark used in this portfolio is NIFTY 50 which contains 50 of the largest Indian companies listed on National Stock Exchange(NSE).
Figure 5: Comparison of a client portfolio before and after rebalancing
We also created a test portfolio by ourselves which is not real rather fabricated. This synthetic portfolio contains of 11 assets from different sectors, Information technology, Health Care, Banking to name a few. We also assumed that there are 100 shares of each of these assets in this Test Portfolio for simplicity. We used various benchmarks to see the efficiency of our rebalancing technique. Fig. 6, 7, 8, shows the the impact of our rebalancing methodology with NIFTY 50, 100, 500 benchmarks respectively. Where as Fig. 9, 10, shows the performance with BSE 100, 500 benchmarks respectively.
Figure 6: Comparison of a test portfolio before and after rebalancing with NIFTY 50 benchmark
Figure 7: Comparison of a test portfolio before and after rebalancing with NIFTY 100 benchmark
Figure 8: Comparison of a test portfolio before and after rebalancing with NIFTY 500 benchmark
Figure 9: Comparison of a test portfolio before and after rebalancing with BSE 100 benchmark
Figure 10: Comparison of a test portfolio before and after rebalancing with BSE 500 benchmark